OpenStudy (anonymous):

Please help! I dont understand this problem! Use summation notation to express the sum.

OpenStudy (anonymous):

\[3-\frac{ 3 }{ 5 }+\frac{ 3 }{ 25 }-\frac{ 3 }{ 125 }+...-\frac{ 3 }{ 390625 }\]